Required information

The Foundational 15 [LO10-1, LO10-2]

[The following information applies to the questions displayed below.]

Westerville Company reported the following results from last year’s operations:

Sales $ 1,000,000
Variable expenses 300,000
Contribution margin 700,000
Fixed expenses 500,000
Net operating income $ 200,000
Average operating assets $ 625,000


At the beginning of this year, the company has a $120,000 investment opportunity with the following cost and revenue characteristics:

Sales $ 200,000
Contribution margin ratio 60 % of sales
Fixed expenses $ 90,000

The company’s minimum required rate of return is 15%.

Foundational 10-1

Required:

5. What is the turnover related to this year’s investment opportunity? (Round your answer to 2 decimal places.)

6. What is the ROI related to this year’s investment opportunity? (Do not round intermediate calculations.Round your answer to the nearest whole percent.)

7. If the company pursues the investment opportunity and otherwise performs the same as last year, what margin will it earn this year? (Round your percentage answer to 1 decimal place (i.e. 0.1234 should be entered as 12.3).)

8. If the company pursues the investment opportunity and otherwise performs the same as last year, what turnover will it earn this year? (Round your answer to 2 decimal places.)

9. If the company pursues the investment opportunity and otherwise performs the same as last year, what ROI will it earn this year? (Do not round intermediate calculations. Round your percentage answer to 1 decimal place (i.e. 0.1234 should be entered as 12.3).)

10-a. If Westerville’s chief executive officer will earn a bonus only if her ROI from this year exceeds her ROI from last year, would she pursue the investment opportunity?

  • Yes

  • No

10-b. Would the owners of the company want her to pursue the investment opportunity?

  • Yes

  • No

11. What is last year’s residual income?

12. What is the residual income of this year’s investment opportunity?

13. If the company pursues the investment opportunity and otherwise performs the same as last year, what residual income will it earn this year?

14. If Westerville’s chief executive officer will earn a bonus only if her residual income from this year exceeds her residual income from last year, would she pursue the investment opportunity?

  • Yes

  • No

15-a. Assume that the contribution margin ratio of the investment opportunity was 50% instead of 60%. If Westerville’s chief executive officer will earn a bonus only if her residual income from this year exceeds her residual income from last year, would she pursue the investment opportunity?

  • Yes

  • No

15-b. Would the owners of the company want her to pursue the investment opportunity?

  • Yes

  • No
Answer #1

ANSWER

5) Turnover = Sales/Average operating assets = 200,000/120,000 = 1.67 Times

6) ROI : Net income/Average operating assets :

Net income = 200,000*60%-90,000= 30,000

ROI = 30,000/120,000 = 25%

7) Margin = (200,000+30,000)/1,200,000= 19.16%

8) Turnover = 1,200,000/745,000 = 1.61 Times

10)  ROI for the the last year = $200,000 / $625,000 = 32%

As ROI for this year decreases from last year by pursuing the new investment opportunity, therefore Westerville’s chief executive officer will not pursue the new investment opportunity.

11) Last year's residual income = Net operating income - Minimum required return = $200,000 - ($625,000*15%) = $106,250

12) Residual income of this year's investment opportunity = $30,000 - ($120,000*15%) = $12,000